OpenStudy (anonymous):

A sample of 145 firecrackers contain 4 duds. How many duds would a box of 1740 firecrackers contain?

OpenStudy (anonymous):

curry are you making these answers up off the top of your head?

OpenStudy (curry):

me?

OpenStudy (curry):

y do u ask that?

OpenStudy (anonymous):

because the last three answer you posted are random

OpenStudy (anonymous):

\[4\times1740\div145\] which equals 48.

OpenStudy (curry):

really?

OpenStudy (anonymous):

yes really

jimthompson5910 (jim_thompson5910):

\[\large \frac{4}{145}=\frac{x}{1740}\] \[\large 4*1740=145x\] \[\large 6960=145x\] \[\large \frac{6960}{145}=x\] \[\large 48=x\] \[\large x=48\] So there would be 48 duds.